Review I Constants Periodic Table Part A The spectral lines of two stars in a particular eclipsing binary system shift back and forth with a period of 10 months. The lines of both stars shift by equal amounts, and the amount of the Doppler shift indicates that each star has an orbital speed of 7.0x104 m/s relative to the other What are the masses of the two stars? Assume that each of the two stars traces a circular orbit around their center of mass. (Hint: See Mathematical Insight Measuring Stellar Masses.) Express your answer using two significant figures. 0図? MSun Submit Request Answer
Answer #1

Given,

The period, T = 10 months

= (10 * 30 * 24 * 60 * 60) sec = 2.6 * 107 sec

Speed of each star, v = 7 * 104 m/s

We know, v = 2\pir / T

=> r = vT / 2\pi

r = (7 * 104 * 2.6 * 107) / 2\pi

r = 18.2 * 1011 / 2\pi

r = 2.897 * 1011 m

By using Kepler's law,

T2 = r3 / (m1 + m2)

T = 10 months

T = (10 /12) years = 0.833 Years

m1 + m2 =  r3 / T2

= (2.897 * 1011)3 / (0.833)2

= 35.04 * 1033  kg

From Doppler shifts, the massive star of mass m2 is twice as m1.

m1 + 2m1 = 35.04 * 1033 kg

3 m1 = 35.04 * 1033 kg

m1 = 11.68 * 1033 kg = 5.872 * 1063 solar mass

So, m2 = 2 * m1

= 2 * 11.68 * 1033

m2 = 23.36 * 1033 kg = 1.174 * 1062 solar mass